Circular No. 6584

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 P. J. Groot and T. J. Galama, University of Amsterdam (UoA); J. van

Paradijs, UoA and University of Alabama, Huntsville; R. Strom, Netherlands Foundation for Research in Astronomy and UoA; J. Telting and R. G. M. Rutten, Isaac Newton Group of Telescopes, La Palma; M. Pettini and N. Tanvir, Institute of Astronomy, Cambridge; R. Naber, University of Groningen; C. Kouveliotou, Universities Space Research Association; J. in ’t Zand and J. Heise, Space Research Organization of the Netherlands, Utrecht; E. Costa, M. Feroci, and L. Piro, Instituto di Astrofisica Spaziale, CNR; F. Frontera and G. Zavattini, Universitá di Ferrara; and L. Nicastro and E. Palazzi, Instituto Tecnologia e Studie Radiazione Extraterrestri, CNR, report: “Comparison of the V - and I-band images of the error box of GRB 970228 (IAUC 6572), taken on Mar. 1.0 UT (IAUC 6574) and on Mar. 8.9 with the William Herschel Telescope (I band, 900 s) and the Isaac Newton Telescope (V band, 2500 s), reveals one object with a large brightness variation in both bands. The variable had V = 21.3 and I = 20.6 on Mar. 1.0 and V > 23.6 and I > 22.2 on Mar. 8.9. Its position is α = 5^{h}01^{m}46^{s}.70, δ = +11^{o}46′53′′.0 (equinox 2000.0; estimated accuracy 1′′). Approximately 4′′ away (at position end figures 46^{s}.73, 53′′.0), we find another object, with V = 23.1 and I = 20.5 on Mar. 8.9, whose brightness on Mar. 1.0 (measured somewhat less accurately because of the nearby variable) differed by < 0.2 mag from these values. The point-spread function of the variable, and that of the nearby constant object, is consistent with those of stars in the images; the two objects are therefore extended by less than ∼ 1′′.5. Radio observations at 5 GHz obtained with the Westerbork Synthesis Radio Telescope on Mar. 1.75 and 2.75 (12 hr each) show no radio source at the position of the variable above 0.35 mJy (3σ). The variable is located within the intersection of the BeppoSAX WFC error box (IAUC 6572), the BeppoSAX LECS error box (IAUC 6576), and the Ulysses/BeppoSAX annulus (IAUC 6578); it is more than 1′ away from the radio source reported on IAUC 6576. Its position and rapid decline contemporaneous with that of the BeppoSAX LECS x-ray transient (IAUC 6576) indicate that the two are related. If the x-ray transient is related to GRB 970228, we have detected the first transient optical signal related to a γ-ray burst.”

Corrigendum. On IAUC 6582, Bloom et al., line 6, for We find V = 21.4, read We find V = 22.4,

1997 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green